# Batch-ICL: Effective, Efficient, and Order-Agnostic In-Context Learning

### v1

@@ 4.2 Experimental Results
We present the experimental results in Table 1. Overall, Batch-ICL demonstrates a substantial improvement over the standard $N$-shot ICL in various tasks and across different model sizes. To determine Batch-ICL’s maximum potential, we identified the optimal $k$ for each test sample and have included these findings in the “Best” column. These results show that the performance discrepancy between our chosen $k$ and the theoretical upper limit of performance is negligible. Furthermore, a larger batch size used during meta-optimization results in reduced noise and randomness in the meta-gradients. Consequently, we observe Batch-ICL exhibits markedly more stable performance across different demonstrations, as evidenced by the significantly reduced variation [^1] An exception occurs with Llama-2-7B in the SST-2 task, where Batch-ICL increases variance. This happens because, in this binary classification task, the “$N$-shot” is close to stable random guessing. in repeated experiments. The improved accuracy and reduced variance serve as initial validation for the effectiveness of Batch-ICL, and also support our explanation in Section 3.1.
@@ 5 Method Analysis
In this section, we delve into various aspects of Batch-ICL, including the time cost, the impact of shot number ($N$), aggregation layer index ($k$), the order of ICL examples, and the optimization “epochs.” This exploration aims to provide a deeper insight into both the ICL task itself and the efficacy of Batch-ICL.
5.1 Discussion on the Efficiency
We analyze the efficiency of Batch-ICL in comparison to the standard $N$-shot ICL approach. Our main focus lies on addressing the computational overhead linked to the attention mechanism in Transformer models. This part is a critical efficiency bottleneck for Transformers, showcasing a time complexity of $O(n^{2})$ when processing texts of length $n$.
@@ 5.2 The Effect of $N$
We first delve into an exploration of the impact of $N$, which plays a crucial role in task settings. The modification of $N$ poses challenges to the robustness in various practical applications. As illustrated in Figure 2, our Batch-ICL consistently surpasses the standard $N$-shot accuracy for different values of $N$. Also, it is important to take note of the gap that exists between the performance of Batch-ICL and its upper limit referred to as “Best.” This gap is minor and diminishes as the value of $N$ increases, implying improved robustness and performance.
Figure 2: Performance dynamics across various $N$.
5.3 The Effect of $k$
@@
We perform a thorough analysis of the impact of $k$. We examined the effects of $k$ by varying its value across the full range of layers, monitoring the performance dynamics as aggregated meta-gradients are implemented at various levels. For each value of $k$, we carried out experiments using the Llama-2-13B model on the SST-2 dataset (Socher et al., 2013), with $N$ values ranging from 2 to 6.
Figure 3 demonstrates that with a small $k$, Batch-ICL does not work. It suggests that the initial, or shallow, layers of LLMs play a crucial role in establishing semantic foundations for the subsequent, deeper layers. This observation aligns with previous studies (Wang et al., 2023; Todd et al., 2023; Hendel et al., 2023).
As the forward computation goes on, we observe an abrupt increase in performance. This improvement plateaus rapidly and does not diminish thereafter. This pattern differs from the findings reported in (Hendel et al., 2023; Todd et al., 2023), where researchers found that a deeper representation of an ICL task diminishes the task-related information it contains. Specifically, when the representation is extracted and used in a zero-shot forward computation, the accuracy tends to converge to zero as the layer depth increases. The discrepancy between ours and theirs findings might be due to subtle differences in input configuration. In our approach, each 1-shot forward computation uses a true query. In contrast, previous studies used a pseudo query, which can lead to significant deviations in deeper layers because most information gathers at the last position (Wang et al., 2023).
@@
Llama-2-7B $60.0$ $90.0$ $65.8$ $63.4$
Llama-2-13B $80.0$ $86.7$ $70.4$ $60.8$
Table 2: The percentage of example orders that Batch-ICL outperforms among all possible example permutations.
@@
1-epoch 2-epoch 3-epoch 4-epoch 1-epoch 2-epoch 3-epoch 4-epoch
2 $57.9$   $58.4$ $58.8$ 59.0    $69.8$    70.0 70.0 70.0
3 $60.3$   $63.4$ $64.8$ 67.0    $71.2$    71.6 71.6 $71.3$
4 $58.3$   $60.2$ $60.7$ 62.2    $69.4$    70.1 $70.0$ $69.7$
5 $57.5$   $59.5$ $60.1$ 60.4    $71.7$    71.8 $71.7$ $71.3$
6 $61.0$   $62.7$ $63.2$ 64.6    $76.2$    76.3 76.3 $75.8$
Table 3: Multiple epochs prove to enhance accuracy, irrespective of model size and shots. Note that, in this study, we do not determine $k$ based on the validation set but rather employ the optimal value for each individual sample.
@@ 5.4 The Effect of ICL Example Order
We now delve into examining the order of ICL examples and its impact on performance.
We meticulously examine all $N!$ possible permutations of ICL examples in a $N$-shot SST-2 (Socher et al., 2013) task. In Figure 4, the “Average” result is the average accuracy of all permutations. “Best Order” denotes the highest accuracy, contrasting with the “Worst Order” which represents the lowest.
@@
Additionally, we performed a statistical analysis on the percentage of permutations that Batch-ICL outperforms. It is obvious that Batch-ICL exceeds the majority of permutations in every comparison. This is particularly noticeable for smaller values of $N$ or with larger models. Overall, this study confirms our motivation, demonstrating that Batch-ICL effectively alleviates concerns related to ICL example orders, often leading to a satisfactory solution.
@@ 5.5 The Effect of Epochs
In Table 3, we present the results of adding more epochs to Batch-ICL on SST-2 (Socher et al., 2013). This table illustrates the model’s performance across various numbers of $N$ and a range of epochs. Our findings indicate that, regardless of the value of $N$ or the size of the model, there is an improvement in the model’s performance over several epochs, culminating in a plateau.
We observed a more pronounced effect in the 7B model compared to the 13B model. The 7B model consistently shows enhanced performance with an increasing number of epochs, peaking at 4 epochs. On the other hand, the 13B model reaches the plateau earlier. This could be due to the larger capacity of the 13B model, which enables it to more effectively capture linguistic subtleties and contextual details with fewer implicit interactions among ICL examples (see Eq.10).
@@ 6 Related Works
The theoretical foundation of our work, including the dual form between linear attention and linear layers optimized by gradient descent, and the interpretation of In-Context Learning (ICL) from a meta-optimization perspective, is grounded in various studies  (Aizerman et al., 1964; Irie et al., 2022; Dai et al., 2023; Ren and Liu, 2023). Additionally, our research is related to the findings of Hendel et al. (2023) and Todd et al. (2023). These works discovered that in few-shot learning tasks, the representation in deep layers carries the task information. Hendel et al. (2023) focused on identifying and aggregating the outputs of induction heads (Olsson et al., 2022). Their work emphasized the content analysis of aggregated hidden states. In contrast, our approach does not differentiate between attention heads, leading to a simpler yet effective method for enhancing the ICL performance. When compared to Hendel et al. (2023), our method demonstrates greater efficacy and surpasses the original few-shot learning performance by using zero-shot learning, an achievement not realized by Hendel et al. (2023). Theoretically, our work offers an explanation for the “training set compression” proposed by Hendel et al. (2023) and elucidates why averaged attention activation, as investigated in (Todd et al., 2023), conveys ICL task information.
@@ 7 Conclusion
In this paper, from a meta-optimization perspective, we provide an explanation for ICL example orders’ impact on performance. Building on our insights, we introduce Batch-ICL, an efficient and effective algorithm for ICL inference. Batch-ICL processes ICL examples in batches, aggregating their meta-gradients. These aggregated meta-gradients are then applied to a zero-shot forward computation for final predictions. Owing to its batch processing approach, Batch-ICL remains agnostic to the order of ICL examples, surpassing the average performance of all order permutations across various tasks. We further enhance Batch-ICL by developing multi-epoch variants that implicitly enumerate permutations of ICL examples. This enhancement fosters better interaction between inputs, thereby further improving the method.
Acknowledgement
This work was supported by Ant Group.

### v3

@@ 4.2 Experimental Results
We present the experimental results on classification tasks in Table 1. Overall, Batch-ICL demonstrates a substantial improvement over both the standard $N$-shot ICL and PCW in various tasks and across different model sizes. Across four out of the five datasets analyzed, Batch-ICL surpasses the performance of F-Ordered, with marginal differences observed on the AGnews dataset.
To determine Batch-ICL’s maximum potential, we identified the optimal $k$ for each test sample and have included these findings in the “Best” column. These results show that the performance discrepancy between our chosen $k$ and the theoretical upper limit of performance is negligible. Furthermore, because Batch-ICL employs a larger batch size during meta-optimization, which diminishes noise and randomness in the meta-gradients, we observe Batch-ICL exhibits markedly more stable performance than standard ICL across different demonstrations, as evidenced by the significantly reduced variation [^2] An exception occurs with Llama-2-7B in the SST-2 task, where Batch-ICL increases variance. This happens because, in this binary classification task, the “$N$-shot” is close to stable random guessing. in repeated experiments. These results serve as initial validation for the effectiveness of Batch-ICL, and also support our explanation in Section 3.2.
Table 2 illustrates the performance comparison for the machine translation task. Batch-ICL outperforms the standard ICL in efficiency. It achieves this while preserving the BLEU score, experiencing only a minimal decrease of 0.7% for the 7B model and 0.3% for the 13B model, respectively.
To showcase the robustness and versatility of Batch-ICL, we have implemented it across different LLMs, including OPT-6.7B Zhang et al. (2022) and Falcon-7B Almazrouei et al. (2023), using SST-2 as a benchmark task. Table 3 illustrates the effectiveness of Batch-ICL across various models, demonstrating its universality.
$N$-shot Batch-ICL
OPT-6.7B $38.4$   $42.5$
Falcon-7B $51.2$   $52.0$
Table 3: Experimental results of SST-2 on a broader array of LLMs.
@@ 5 Method Analysis
We delve into various aspects of Batch-ICL, including the impact of shot number ($N$), aggregation layer index ($k$), the impact of the order of ICL examples, and the optimization “epochs.” This exploration aims to provide a deeper insight into both the efficacy of Batch-ICL and the ICL itself.
@@ 5.1 The Effect of $N$
(a) SST-2
(b) RTE
Figure 2: Performance dynamics across various $N$ on SST-2 and RTE.
Different tasks typically demand varying values of $N$. The effectiveness of Batch-ICL across different $N$ represents its robustness in various practical applications. As illustrated in Figure 2, our Batch-ICL consistently surpasses the standard $N$-shot accuracy for different values of $N$. Also, the gap between the performance of Batch-ICL and its upper limit referred to as “Best” is minor and diminishes as the value of $N$ increases.
$N=4$ $N=10$ $N=20$ $N=70$
7B 29.3 36.24 39.43 33.3
13B 31.4 32.67 33.70 39.1
Table 4: Results from Batch-ICL on the AGNews task with varying values of $N$. Even when $N=70$, demonstration examples surpass the model’s maximum context length, yet it still achieves additional improvements.
@@
Llama-2-7B $60.0$ $90.0$ $65.4$ $63.7$
Llama-2-13B $80.0$ $88.3$ $71.3$ $60.5$
Table 5: The percentage of permutations that Batch-ICL outperforms.
Due to the parallel processing of ICL examples, as long as each individual example doesn’t exceed the maximum context length, there’s no limit to the number of examples Batch-ICL can handle. In Table 4, for the AGNews task, as $N$ increases from $4$ to $70$, we found the performance first hits a peak and then drops on Llama-2-7B; By contrast, Batch-ICL keeps bringing benefits with a larger $N$ on Llama-2-13B. We assume the drop in performance at $N=70$ on Llama-2-7B may stem from the fact that 7B model has already exploited its ability, while 13B is stronger at learning from more shots. Notably, even when we include $N=70$ examples, averaging a total of 5,447 tokens, which significantly exceeds Llama-2’s context limit of 4,096, Batch-ICL continues to demonstrate additional enhancements.
These analyses demonstrate that Batch-ICL is robust to $N$ and can more thoroughly exploits LLMs’ in-context learning capacity.
@@
Epoch 1 2 3 4 Epoch 1 2 3 4
2 SST-2 $57.97$ $58.78$ $58.91$ 59.48 SST-2 $70.09$ 70.30 $69.79$ $69.83$
RTE $67.7$ 67.8 $67.7$ $67.9$ RTE $79.7$ 80.0 $79.5$ $79.7$
3 SST-2 $60.49$ $64.27$ $65.11$ 67.77 SST-2 71.58 $71.35$ $71.39$ $71.40$
RTE $68.2$ 68.3 $68.3$ $68.3$ RTE $79.9$ 80.1 $80.0$ $80.0$
4 SST-2 $58.84$ $60.52$ $60.60$ 62.10 SST-2 70.19 $69.90$ $69.93$ $69.83$
RTE $67.6$ 67.9 $67.8$ $67.7$ RTE $80.9$ $80.9$ $81.1$ 81.3
5 SST-2 $57.80$ $59.96$ $60.32$ 61.39 SST-2 $71.80$ 71.82 $71.37$ $71.26$
RTE $68.2$ $68.3$ 68.4 $68.1$ RTE $81.4$ 81.6 $81.4$ $81.4$
6 SST-2 $61.49$ $63.08$ $63.32$ 64.57 SST-2 $76.52$ 76.55 $76.26$ $75.84$
RTE $67.6$ $67.7$ $67.8$ 67.9 RTE $81.8$ 82.0 $82.0$ $82.0$
Table 6: Multiple epochs prove to enhance accuracy on SST-2 and RTE, irrespective of model size and shots. Note that, in this study, we do not determine $k$ based on the validation set but rather employ the optimal value for each individual sample.
5.2 The Effect of $k$
We examined the effects of $k$ by varying its value across the full range of layers, monitoring the performance dynamics as aggregated meta-gradients are implemented at various levels. For each value of $k$, we carried out experiments using the Llama-2-13B model on the SST-2 dataset (Socher et al., 2013), with $N$ values ranging from 2 to 6.
Figure 3 demonstrates that with a small $k$, Batch-ICL does not work. It suggests that the initial, or shallow, layers of LLMs play a crucial role in establishing semantic foundations for the subsequent, deeper layers. This observation aligns with previous studies (Wang et al., 2023b; Todd et al., 2023; Hendel et al., 2023).
As the forward computation goes on, we observe an abrupt increase in performance. This improvement plateaus rapidly and does not diminish thereafter. This pattern differs from the findings reported in (Hendel et al., 2023; Todd et al., 2023), where researchers found that a deeper representation of an ICL task diminishes the task-related information it contains. Specifically, when the representation is extracted and used in a zero-shot forward computation, the accuracy tends to converge to zero as the layer depth increases. The discrepancy between ours and theirs findings might be due to subtle differences in input configuration. In our approach, each 1-shot forward computation uses a true query. In contrast, previous studies used a pseudo query, which can lead to significant deviations in deeper layers because most information gathers at the last position (Wang et al., 2023b).
@@ 5.3 The Effect of ICL Example Order
We meticulously examine all $N!$ possible permutations of ICL examples in a $N$-shot SST-2 (Socher et al., 2013) task. Due to the prohibitively high cost associated with enumerating the permutations, we restrict the value of $N$ within the range of 2 to 5. In Figure 4, the “Average” result is the average accuracy of all permutations. “Best Order” denotes the highest accuracy, contrasting with the “Worst Order” which represents the lowest.
@@
Additionally, we performed a statistical analysis on the percentage of permutations that Batch-ICL outperforms. In Table 5, it is obvious that Batch-ICL exceeds the majority of permutations for every $N$. This is particularly noticeable for a smaller $N$ and a larger model. Overall, this study confirms our motivation, demonstrating that Batch-ICL effectively alleviates concerns related to ICL example orders, leading to a satisfactory solution.
@@ 5.4 The Effect of Epochs
In Table 6, we present the results of adding more epochs to Batch-ICL on SST-2 (Socher et al., 2013) and RTE Dagan et al. (2006); Wang et al. (2019). This table illustrates the model’s performance across various numbers of $N$ and a range of epochs. Our findings indicate that, across the majority of $N$ and model sizes, there is an improvement in the model’s performance over several epochs. Specifically, in the RTE task, we observe a performance plateau typically achieved within 2 epochs. Conversely, in the SST-2 task, we extend our investigation across more epochs, typically reaching a plateau after approximately 10 to 20 epochs, as illustrated in Table 7.
Epoch=4 Epoch=10 Epoch=20
7B 62.10 67.84 67.40
13B 69.83 76.30 72.38
Table 7: Performance of Batch-ICL with more epochs.
Overall, we observed a more pronounced effect in the 7B model compared to the 13B model. The 7B model consistently shows enhanced performance with an increasing number of epochs, peaking at 4 epochs. On the other hand, the 13B model reaches the plateau earlier. This could be due to the larger capacity of the 13B model, which enables it to more effectively capture linguistic subtleties and contextual details with fewer implicit interactions among ICL examples (see Eq.10).
@@ 6 Related Works
In addition to PCW we compared in §4, our study is also related to the findings of Hendel et al. (2023) and Todd et al. (2023). These works discovered that in few-shot learning tasks, the representation in deep layers carries the task information. Hendel et al. (2023) focused on identifying and aggregating the outputs of induction heads (Olsson et al., 2022). They emphasized the content analysis of aggregated hidden states. In contrast, we focus on a simpler yet effective method for enhancing the ICL and do not differentiate specific heads. When compared to Hendel et al. (2023), our method demonstrates greater efficacy and surpasses the original few-shot learning performance by using zero-shot learning, an achievement not realized by Hendel et al. (2023). Theoretically, our work offers an explanation for the “training set compression” proposed by Hendel et al. (2023) and elucidates why averaged attention activation conveys ICL task information (Todd et al., 2023).
@@ 7 Conclusion
From a meta-optimization perspective, we explain ICL example orders’ impact on performance. Building on our insights, we introduce Batch-ICL, an efficient and effective algorithm for ICL inference. Batch-ICL processes ICL examples in batches, aggregating their meta-gradients. Aggregated meta-gradients are then applied to a zero-shot forward computation for final predictions. Due to the batch processing, Batch-ICL is agnostic to the order of ICL examples, surpassing the average performance of all order permutations across various tasks, and supports much more examples. We expand Batch-ICL by developing multi-epoch variants that implicitly enumerate permutations of ICL examples, which fosters better interaction between inputs and further improves our method.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (NSFC Grant No. 62122089), Beijing Outstanding Young Scientist Program NO. BJJWZYJH012019100020098, and Intelligent Social Governance Platform, Major Innovation & Planning Interdisciplinary Platform for the "Double-First Class" Initiative, Renmin University of China, the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China. This work was supported by Ant Group Research Fund. Ang Lv is supported by the Outstanding Innovative Talents Cultivation Funded Programs 2023 of Renmin University of China.
Limitations
The theoretical foundation of our work is grounded in various studies  (Aizerman et al., 1964; Irie et al., 2022; Dai et al., 2023) which takes the attention in Transformers as a linear attention. Some research Ren and Liu (2023) suggests that even without this simplification, the conclusions and insights of these studies remain valid. Nevertheless, for the sake of clarity in presentation and ease of comprehension, we adhere to the linear simplification. Additionally, in the section where we explore multiple “epochs,” we simplify the feed-forward layer as a linear transformation. This simplification is widely adopted in many works on interpretability Olsson et al. (2022); Wang et al. (2023a); Yu et al. (2023); Wang et al. (2023c) due to the considerable challenges associated with analyzing nonlinear MLP.
The potential risks of our study are similar to those of other works involving LLMs, as they can sometimes generate toxic responses.